OpenStudy (anonymous):

To test that the solution to the equation below is z = –2, which of the following calculations should be performed?

OpenStudy (anonymous):

2 = 2z + 16 A. 12=2*(-2)+16 B. 12 + (–2) = 2z + 16 + (–2) C. 12 – (–2) = 2z + 16 – (–2) D. 12*(-2)=2z*(-2)+16*(-2)

OpenStudy (cwrw238):

is your equation correct shouldn't it be 12 = 2z + 16 ?

OpenStudy (anonymous):

oh yes you are right i forgot the 1 my bad

OpenStudy (cwrw238):

work out A is right hand side = 12?